Check whether the following are quadratic equations: (i) (x + 1)^2 = 2(x – 3)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
(x + 1)^2 = 2(x – 3)
By using the formula for (a+b)^2= a^2+2ab+b^2
⇒ x^2 + 2x + 1 = 2x – 6
⇒ x^2 + 7 = 0
Since the above equation is in the form of ax2 + bx + c = 0.
Therefore, the given equation is quadratic equation.

Thanks